OpenStudy (anonymous):

need help, can't get the right answer complex fraction 2x^-2 + y^-2/x^-1 - y^-1

OpenStudy (espex):

Is this the question? \[\frac{2x^{-2} + y^{-2}}{x^{-1} - y^{-1}}\]

OpenStudy (anonymous):

yes that's it

OpenStudy (anonymous):

\[\{2x{}^{\wedge}-2+y{}^{\wedge}-2\}/(x{}^{\wedge}-1-y{}^{\wedge}-1)=\frac{\frac{2}{x^2}+\frac{1}{y^2}}{\frac{1}{x}-\frac{1}{y}}=\frac{3}{y-x}+\frac{2}{x}-\frac{1}{y} \]

OpenStudy (anonymous):

the first part is right, thanks, i got it